4) You are considering investing in a local business that modifies equipment & sells it to the public at a much higher price.Please show me your clear work. thank you so much!!
Answer #1

Selling price per unit = $25,000 p.u.

Variable Cost per unit (labor & materials) = $10,000

Contribution per unit = 25,000 – 10,000 = $15,000

Annual Contribution = 15,000*4 units = $60,000

.

Cost of Fixed Assets = $60,000

Project Life = 4 years

Annual Depreciation = Cost of New Sausage system/ Project Life

= 60,000/4

= $15,000

.

.

Profit on sale of Sausage system at the end of project = Sales Price – Book Value

= 10,000 – 0

= $10,000

After Tax Salvage Value of Sausage system at the end of project = Profit on sale of Sausage system at the end of project*(1-Capital gain tax rate)

= 10.000*(1-0.15)

= 8,500

.

.Calculation of NPV of the project:

Year Inititial Investment (1) Working Capital (2) Contribution from Sales (3) Lease Expenses (4) Depreciation (5) Net Profit Before Tax = (6) = (3)-(4)-(5) Tax @ 34% (7) = (6)*0.34 Net Profit after Tax (8) = (6)-(7) Cash Flow from Operations (9) = (8)+(5) After Tax Salavge Value (10) Net Cash Flow (11) = (1)+(2)+(9)+(10) DF Working Discounting Factor @9% (12) Present Value (13) = (11)*(12)
0                                (60,000)                         (10,000)                                                                          -                                                                            -                                -                                                                                   -                                                      -                                                                                  -                                                                                  -                                                       -                                                                      (70,000)                      1                                                       1                                        (70,000.00)
1                                            -                                       -                                                                  60,000                                                                15,000                     15,000                                                                        30,000                                           10,200                                                                      19,800                                                                      34,800                                                     -                                                                        34,800 1/1.14^1 0.877192982                                          30,526.32
2                                            -                                       -                                                                  60,000                                                                15,000                     15,000                                                                        30,000                                           10,200                                                                      19,800                                                                      34,800                                                     -                                                                        34,800 1/1.14^2 0.769467528                                          26,777.47
3                                            -                                       -                                                                  60,000                                                                15,000                     15,000                                                                        30,000                                           10,200                                                                      19,800                                                                      34,800                                                     -                                                                        34,800 1/1.14^3 0.674971516                                          23,489.01
4                                            -                             10,000                                                                60,000                                                                15,000                     15,000                                                                        30,000                                           10,200                                                                      19,800                                                                      34,800                                              8,500                                                                      53,300 1/1.14^4 0.592080277                                          31,557.88
NPV:                                          42,350.67

.

Therefore, maximum price we can pay for the project is its Net Present Value i.e.$42,350.67